OpenStudy (henryarias14):

FINDING THE DERIVATIVE F(x)=|1/x| that's an absolute value. I have a good knowledge of how to find derivatives but when I saw the absolute value I blanked out. I'd appreciate any help, thank you

OpenStudy (solomonzelman):

In general, the derivative of an absolute value function \(y=|f(x)|\) is given as \(y'=\displaystyle \frac{f(x)}{|f(x)|}\times f'(x)\).

OpenStudy (henryarias14):

So it would be 1 • 1/x ?

OpenStudy (solomonzelman):

\(\displaystyle \frac{1/x}{|1/x|}\times \left(\frac{d}{dx} (1/x)\right)\)

OpenStudy (solomonzelman):

if you go according to that.

OpenStudy (solomonzelman):

You could simplify it in several ways, but what you did is not correct.

OpenStudy (henryarias14):

I understand but what happens with the absolute value in the denominator?

OpenStudy (3mar):

Thank you, SolomonZelman It is really helpful for me!

OpenStudy (henryarias14):

Could you simplify it ?

OpenStudy (henryarias14):

I know it's 1/x / |1/x • 1/x^2

OpenStudy (solomonzelman):

actually it is -1/x^2.

OpenStudy (3mar):

I think it will be \[y'=\frac{ -1 }{ x^3|\frac{ 1 }{ x }| }\]

OpenStudy (henryarias14):

Yeah that's what I meant

OpenStudy (solomonzelman):

\(\displaystyle \frac{1/x}{|1/x|}\times \left(\frac{d}{dx} (1/x)\right)=\frac{1/x}{|1/x|}\times \left(\frac{-1}{x^2}\right)=\) \(\displaystyle (1/x)\times (|x|/1)\times\frac{-1}{x^2} =-\frac{|x|}{x^3}=-\frac{|x|}{|x^2|x}=\frac{-1}{|x|x}\).

OpenStudy (3mar):

Great!

OpenStudy (solomonzelman):

I will give you the derivation of it. ------------------------------ PROOF (not quite rigorous, but....) Let \(f\) be a function of \(x\), and let \(y=|f|\). Assume \(f\) is differentiable, and assume \(x\in \mathbb{R}\). (You'll see why the 2nd assumption is there.) Definition of Absolute Value (for real numbers): \(z=\sqrt{z^2}\). Rewriting the function this way gives \(y=\sqrt{f^2}\). (Don't forget, \(f\) is a function of \(x\).) Applying the rules of differentiation: \(\displaystyle y'=\frac{1}{2\sqrt{f^2}}\times \left[\frac{d}{dx}(f^2)\right]=\frac{1}{2\sqrt{f^2}}\times \left[2f\times f'\right]=\frac{f\times f'}{\sqrt{f^2}}\). applying the definition of the absolute value, back, \(\displaystyle y'=\frac{f\times f'}{|f|}\).

OpenStudy (solomonzelman):

This is why / Thus .... \(\displaystyle\frac{d}{dx}|f(x)|= \frac{f(x)}{|f(x)|}\times f'(x)\).

OpenStudy (henryarias14):

One question though. Why did you make the x^3 into |x^2|x

OpenStudy (henryarias14):

And how

OpenStudy (3mar):

nice ques!

OpenStudy (solomonzelman):

\(x^2=|x^2|\) for any real x. (Right?) Hence, \(x^3=x^2\times x=|x^2|x\).

OpenStudy (solomonzelman):

or, do you mean by "why" - "to accomplish what did I do this" ?

OpenStudy (henryarias14):

I see. But why not make it into |x|x^2 ?

OpenStudy (henryarias14):

Nono i know why you did it

OpenStudy (solomonzelman):

if \(x\) is negative, then \(|x|x^2\) is positive and \(x^3\) is negative. They are (equivalent in magnitude, but) not equivalent in sign.

OpenStudy (solomonzelman):

questions?

OpenStudy (henryarias14):

No i got it now, I was trying to work to out thanks!

OpenStudy (henryarias14):

Helped a lot

OpenStudy (solomonzelman):

Tnx for the feed:)